A Ring of radius R is first rotated with an angular velocity ∆ and then carefully placed on a horizontal rough surface. The coefficient of friction between the surface and the ring is ¶. Time after which its angular speed is reduced to half is
1
Expert's answer
2019-04-24T10:10:36-0400

Let "\mu" be the coefficient of friction, "\omega" be the angular speed and "\alpha" the angular acceleration. Consider a torque of force of friction:

"\tau_\text{f}=\mu mg\cdot R,"

and according to Newton's second law for rotational motion, the net torque is


"\tau_\text{net}=mR^2\alpha=mR^2\cdot \frac{\omega-\omega/2}{t},"

equate them:


"\mu mg\cdot R=mR^2\cdot \frac{\omega}{2t},"

"t=\frac{\omega R}{2\mu g}."